# Almost Complete Intersections: Regularity of powers and Rees algebra

### v1

@@ Abstract.
An almost complete intersection ideal can be seen as a $d$-sequence ideal with the minimal number of generators being one more than its height. In this paper, we give exact formulas for the regularity of powers of graded almost complete intersection ideals. We also present the minimal bigraded free resolutions of Rees algebras associated with linear type ideals having one generator more than their grade and study the properties of Cohen-Macaualyness, Koszulness associated to their diagonals.
@@ Introduction
All rings in this paper are considered to be standard graded $K$-algebras ($K$ being a field), unless otherwise specified. Rees algebras and symmetric algebras are important classes of bigraded algebras which have powers and symmetric powers of ideals respectively as its graded components. Corresponding to a graded ideal, the defining equations of the symmetric algebra can be given by the presentation of the ideal, but it is not easy to obtain the associated free resolution. Whereas in case of Rees algebras, even finding the explicit defining equations are found to be challenging in general. Ideals of linear type are a class of ideals for which the corresponding symmetric algebras and Rees algebras are isomorphic, which is equivalent to the defining equations of Rees algebras being linear in the second set of variables. Complete intersection ideals and $d$-sequence ideals are some classes of ideals of this type. In case of ideals generated by regular sequences, Koszul complexes resolve the ideals and the Eagon-Northcott complex resolves the minimal free resolution of the associated Rees algebra ([11, 12]). In [24], Rees algebras of certain Pfaffians ideals generated by $d$-sequence which come corresponding to alternating matrices are discussed. There are other special cases for which the correponding Rees algebras are studied [11, 33, 34, 28, 14, 26].
@@
Another class of ideals of linear type are almost complete intersections. An ideal $I$ in a standard graded $K$-algebra $A$ is said to be an almost complete intersection if the minimal number of generators of $I$ is one more than the height of $I$ along with the property that for all minimal primes $\mathfrak{p}$ of $I$ in $A$, $I_{\mathfrak{p}}$ is a complete intersection. It is proved in [21] that these ideals are contained in the class of $d$-sequence ideals. Further some study has been done on their regularity [20, 31], combinatorial characterization of graphs which correspond to these ideals ([21]) and their corresponding Rees algebras [13, 21]. These classes of ideals can be seen as losing the complete intersection property by one generator, yet unlike the ideals generated by regular sequences, little is known in generality regarding the homological invariants associated to them.
@@
Minimal free resolutions of ideals are one of the main tools which helps in studying their structure. Finding these resolutions corresponding to arbitrary ideals in general, is a difficult problem. Applying stronger conditions on the ideal helps in a better understanding of the homological invariants associated to them. In Section 3 (Proposition 3.1) we give the minimal bigraded free resolution of the Rees algebra corresponding to an ideal satisfying the following properties: $(1)$ ideal is of linear type, $(2)$ minimal number of generators of the ideal is one more than its grade. This is a generalization of a result in [5] extending it to the case of non-equigenerated ideals. As a consequence, one obtains resolutions of the bigraded algebras associated to almost complete intersection ideals. In particular, it shows that in order to obtain the minimal free resolution of the Rees algebra of almost complete intersection ideals, it suffices to have information about the resolution of $I$.
@@
A significant algebraic invariant which measures the complexity of a module is Castelnuovo-Mumford regularity. It is proved in [10, 22] that for every homogeneous ideal $I$, the regularity of $I^{s}$ is of the form $as+b$ for $s\gg 0$ and non-negative constants $a,b$. While computing the value of $b$ is often considered to be a challenging problem, the value of $a$ is widely understood in literature. In Section 2 (Theorem 2.3), we derive the constant $b$ for powers of almost complete intersections thereby extending the result of Shen and Zhu [31, Theorem 3.1] to include the case when the ideal is non-equigenerated.
@@
A standard graded $K$-algebra is Koszul if the minimal free resolution of the corresponding residue field $K$ has a linear resolution. Some articles that have discussed these properties of the diagonal subalgebras of Rees algebras of ideals are [32, 9, 6, 23, 1] which includes the cases of complete intersections, residual intersections etc. In Section 4, we study the ring-theoretic properties like Cohen-Macaulayness, Koszulness of the diagonals of Rees algebras associated to certain specific classes of ideals of linear type. For a complete intersection ideal $I$ in a polynomial ring minimally generated by $r$ elements of degree $d$, it is proved in [9, Corollary 6.10] that for $c\geq\frac{d(r-1)}{r}$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul and an improved bound of $c\geq\frac{d}{2}$ is given for the case of $r=2$, $d=2$ in [6, Corollary 3.3] and for the general case of $r=2$ and any $d$ in [23, Theorem 3.1]). In Proposition 4.1$(1)$, we prove a similar result for an almost complete intersection ideal $I$ minimally generated by $3$ elements in degree $d$ in a graded polynomial ring, which is, $\mathcal{R}(I)_{\Delta}$ is Koszul for $c\geq\frac{d}{3}$ and $e>0$.
@@ 1. Preliminaries
Following are some notations used throughout the article.
@@
Set $t_{\mathnormal{i}}^{A}(M):=\sup\{\mathnormal{j}:\beta_{\mathnormal{i},\mathnormal{j}}^{A}(M)\neq 0\}$ where $\beta_{i,j}$ is the $(i,j)^{\text{th}}$ Betti number of $M$ and by convention, $t_{\mathnormal{i}}^{A}(M)=-\infty$ if $F_{\mathnormal{i}}=0.$ Castelnuovo-Mumford regularity is then defined as follows:
@@
Castelnouvo-Mumford regularity measures the complexity of the minimal free resolution of a finitely generated $A$-module $M$. We use the notation $\operatorname{reg}(M)$ when the ring $A$ is evident.
@@
[17, Lemma 3.1] Let $0\rightarrow M\rightarrow N\rightarrow P\rightarrow 0$ be a short exact sequence of finitely generated graded $A$-modules. Then, the following holds.
@@
Following are some properties of $d$-sequences and regular sequences which will be used in this article.
@@
    If $\{f_{1},\ldots,f_{n}\}$ forms a $d$-sequence in a ring $A$, then
    1. (a)
      $(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n}^{s})=(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n})$ for all $s\geq 1$.
    2. (b)
      For $s\geq 1$, ${(\langle f_{1},\ldots,f_{i-1}\rangle+I^{s}:f_{i}})=(\langle f_{1},\ldots,f_{i-1}\rangle:f_{i})+I^{s-1}$ for all $1\leq i\leq n$ ([31, Observation 2.4]).
@@
Following are some results related to the Koszulness and Cohen-Macaulayness of the diagonals of bigraded algebras which will be used in the paper.
@@
    For $\Delta=(c,e)$ and a bigraded $K$-algebra $R$ with the $i^{th}$ Betti number corresponding to the shift $(a,b)$ being denoted by $\beta_{i,a,b}$, $R_{\Delta}$ is Koszul if $\max\{\dfrac{a}{c},\dfrac{b}{e}:\beta_{i,a,b}\neq 0\}\leq i+1$ for every $i$ ([9, Theorem 6.2]).
@@ 2. Regularity of powers of Almost Complete Intersections
In this section, we extend their result to give the regularity of powers of almost complete intersections in the non-equigenerated case as well. Our result coincides with the result in [31] when the ideal is equigenerated. Note that an upper bound for the same is given in [20].
@@
Consider the notation as assumptions as in Setup 1. Then,
@@
Here, $\operatorname{reg}\left(\dfrac{A}{J}\right)=\sum_{l=1}^{n-1}d_{l}-(n-1)>\operatorname{reg}\left(\dfrac{A}{I}\right)$. Hence,
@@
As a consequence of Note 1.3, $\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}\right)=\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}\rangle}\right)=\sum_{l=1}^{n-1}d_{l}-n+1-d_{n}.$ Then $\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}(-2d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{J}\right)+d_{n}$.
@@
Since $\{f_{1},\ldots,f_{n}\}$ forms a $d$-sequence with $\{f_{1},\ldots,f_{n-1}\}$ forming a regular sequence, from Note 1.3,
$\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}):f_{i}}\right)(-d_{i})=\operatorname{reg}\left(\dfrac{A}{I}\right)+d_{i}<\sum_{l=1}^{n-1}d_{i}-(n-1)+d_{i}\leq\sum_{l=1}^{n}d_{i}-(n-1)\leq$
$\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\right)+1$. Hence from Lemma $(1)$,
@@
From Note 1.3, one obtains $\operatorname{reg}\left(\dfrac{A}{J:f_{n}^{s+1}}\right)(-sd_{n})=\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)+sd_{n}$. Then from Lemma 1.1$(3)$,
@@
As a consequence of the above results, one can obtain the regularity of powers of almost complete intersections.
@@
Consider the notation and assumptions as in Setup 1. Then for all $s\geq 2$ and $i=0,\ldots,n-1$, the following is true,
@@
We prove the result by descending induction on $i$. Consider the base case of $i=n-1$ of the descending induction in Equation (4). Then clearly, $\langle f_{1},\ldots,f_{(n-1)+1}\rangle+I^{s}=I$. From Note 1.3 and Lemma 2.2,
$\operatorname{reg}\left(\dfrac{A}{(J+I^{s}):f_{n}}(-d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+I^{s-1}}\right)+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}\geq\sum_{l=1}^{n-1}d_{l}-(n-1)+1>\operatorname{reg}\left(\dfrac{A}{I}\right)+1$. Hence $\operatorname{reg}\left(\dfrac{A}{J+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}$ by Lemma $(1)$.
Now assume $i\leq n-2$ and that the result holds for $i+1$. In the short exact sequence (4), by induction hypothesis on $i$, $\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}$. Since $\{f_{1},\ldots,f_{j}\}$ forms a $d$-sequence for $1\leq j\leq n$ and a regular sequence for $11\leq j\leq n-$ 1, from Note 1.3, $\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)=\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{s-1}}\right)$. This by induction hypothesis on $s$ is equal to $\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}$.
Now, the final result is a consequence of the inequality
$\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)+d_{i+1}<\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)+1$
and Lemma $(1)$. ∎
@@
The above theorem fails if $d_{n}$ is not the maximal degree generator of $I$. For example consider $I=\langle f_{1},f_{2},f_{3},f_{4}\rangle=\langle x_{26}x_{45}-x_{25}x_{46}+x_{24}x_{56},x_{16}x_{45}-x_{15}x_{46}+x_{14}x_{56},x_{16}x_{25}x_{34}-x_{15}x_{26}x_{34}-x_{16}x_{24}x_{35}+x_{14}x_{26}x_{35}+x_{15}x_{24}x_{36}-x_{14}x_{25}x_{36},x_{36}x_{45}-x_{35}x_{46}+x_{34}x_{56}\rangle\subset K[x_{14},x_{15},x_{16},x_{24},x_{25},x_{26},x_{34},x_{35},x_{36},x_{45},x_{46},x_{56}]=B$. Then $\{f_{1},f_{2},f_{3},f_{4}\}$ forms a $d$-sequence with $\{f_{1},f_{2},f_{3}\}$ being a regular sequence. But $\operatorname{reg}\left(\dfrac{B}{I^{3}}\right)=8$ which is not equal to $(2+2+2+3)-4+2=7$.
As a consequence of the above results, the following can be observed regarding the regularity of powers of almost complete intersections of grade $3$. To understand the set of generators of these ideals, we look at the notion of Pfaffians.
Given a skew-symmetric matrix $X$, the Pfaffian of $X$ is defined as the square root of the determinant of $X$. For $l\in\mathbb{N}$, let $\operatorname{Pf}_{\bar{l}}(X)$ denotes the Pfaffian of $X$ obtained by removing the $l^{th}$ row and the corresponding column of $X$ which can be seen as an element in $B=K[X]$. In [7], the authors show that almost complete intersections of grade $3$ arise as the ideal generated by certain Pfaffians of skew-symmetric matrices of some particular form (see Setup 3). Following are some properties of these ideals that can be observed.
@@
    $B/I$ is Cohen-Macaulay. Since $I$ is perfect in $B$, the observation follows from [12, Section 19].
@@
    For $t\geq 5$ and $I$ an ideal of type $t-3$,
    1. (a)
      $\text{reg}_{B}(B/I)=t-4$.
    2. (b)
      $(B/I)^{(c)}$ is Koszul for $c\geq\dfrac{t}{3}$.
@@
Assume the notation and hypotheses as in Setup 3. Then,

### v2

@@ Abstract.
An almost complete intersection ideal can be seen as a $d$-sequence ideal with the minimal number of generators being one more than its height. In this paper, we give exact formulas for the regularity of powers of graded almost complete intersection ideals satisfying certain conditions. We also present the minimal bigraded free resolutions of Rees algebras associated with linear type ideals having one generator more than their grade and study the properties of Cohen-Macaulayness, Koszulness associated to their diagonals.
@@ Introduction
All rings in this paper are considered to be standard graded $K$-algebras ($K$ being a field), unless otherwise specified. Rees algebras and symmetric algebras are important classes of bigraded algebras which have powers and symmetric powers of ideals respectively as its graded components. Corresponding to a graded ideal, the defining equations of the symmetric algebra can be given by the presentation of the ideal, but it is not easy to obtain the associated free resolution. Whereas in case of Rees algebras, even finding the explicit defining equations are found to be challenging in general. Ideals of linear type are a class of ideals for which the corresponding symmetric algebras and Rees algebras are isomorphic, which is equivalent to the defining equations of Rees algebras being linear in the second set of variables. Complete intersection ideals and $d$-sequence ideals are some classes of ideals of this type. In case of ideals generated by regular sequences, Koszul complexes resolve the ideals and the Eagon-Northcott complex resolves the minimal free resolution of the associated Rees algebra ([12, 13]). In [25], Rees algebras of certain Pfaffians ideals generated by $d$-sequence which come corresponding to alternating matrices are discussed. There are other special cases for which the corresponding Rees algebras are studied [12, 33, 34, 28, 15, 26].
@@
Another class of ideals of linear type are almost complete intersections. An ideal $I$ in a standard graded $K$-algebra $A$ is said to be an almost complete intersection if the minimal number of generators of $I$ is one more than the height of $I$ along with the property that for all minimal primes $\mathfrak{p}$ of $I$ in $A$, $I_{\mathfrak{p}}$ is a complete intersection. It is proved in [22] that homogeneous almost complete intersection ideals are contained in the class of $d$-sequence ideals (Proposition 1.5). Further some study has been done on their regularity [21, 31], combinatorial characterization of graphs which correspond to these ideals ([22]) and their corresponding Rees algebras [14, 22]. These classes of ideals can be seen as losing the complete intersection property by one generator, yet unlike the ideals generated by regular sequences, little is known in generality regarding the homological invariants associated to them.
@@
Minimal free resolutions of ideals are one of the main tools which helps in studying their structure. Finding these resolutions corresponding to arbitrary ideals in general, is a difficult problem. Applying stronger conditions on the ideal helps in a better understanding of the homological invariants associated to it. In Section 3 (Theorem 3.1) we give the minimal bigraded free resolution of the Rees algebra corresponding to an ideal satisfying the following properties: $(1)$ ideal is of linear type, $(2)$ minimal number of generators of the ideal is one more than its grade. This is a generalization of a result in [5] extending it to the case of non-equigenerated ideals. As a consequence, one obtains resolutions of the bigraded algebras associated to almost complete intersection ideals. In particular, it shows that in order to obtain the minimal free resolution of the Rees algebra of almost complete intersection ideals, it suffices to have information about the resolution of $I$.
@@
A significant algebraic invariant which measures the complexity of a module is the Castelnuovo-Mumford regularity. It is proved in [11, 23] that for every homogeneous ideal $I$, the regularity of $I^{s}$ is of the form $as+b$ for $s\gg 0$ and non-negative constants $a,b$. While computing the value of $b$ is often considered to be a challenging problem, the value of $a$ is widely understood in literature. In Section 3 (Theorem 2.3), we derive the constant $b$ for powers of almost complete intersections thereby extending the result of Shen and Zhu [31, Theorem 3.1] to include the case when the ideal is non-equigenerated.
@@
A standard graded $K$-algebra is Koszul if the minimal free resolution of the corresponding residue field $K$ has a linear resolution. Some articles that have discussed these properties of the diagonal subalgebras of Rees algebras of ideals are [32, 10, 7, 24, 1] which includes the cases of complete intersections, residual intersections etc. In Section 4, we study the ring-theoretic properties like Cohen-Macaulayness, Koszulness of the diagonals of Rees algebras associated to certain specific classes of ideals of linear type. For a complete intersection ideal $I$ in a polynomial ring minimally generated by $r$ elements of degree $d$, it is proved in [10, Corollary 6.10] that for $c\geq\frac{d(r-1)}{r}$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul and an improved bound of $c\geq\frac{d}{2}$ is given for the case of $r=2$, $d=2$ in [7, Corollary 3.3] and for the general case of $r=2$ and any $d$ in [24, Theorem 3.1]). In Proposition 4.1$(1)$, we prove a similar result for an almost complete intersection ideal $I$ minimally generated by $3$ elements in degree $d$ in a graded polynomial ring, which is, $\mathcal{R}(I)_{\Delta}$ is Koszul for $c\geq\frac{d}{3}$ and $e>0$.
@@ 1. Preliminaries
The following are some notations used throughout the article.
@@
Set $t_{\mathnormal{i}}^{A}(M):=\sup\{\mathnormal{j}:\beta_{\mathnormal{i},\mathnormal{j}}^{A}(M)\neq 0\}$ where $\beta_{i,j}$ is the $(i,j)^{\text{th}}$ Betti number of $M$ and by convention, $t_{\mathnormal{i}}^{A}(M)=-\infty$ if $F_{\mathnormal{i}}=0.$ The Castelnuovo-Mumford regularity is then defined as follows:
@@
The Castelnuovo-Mumford regularity measures the complexity of the minimal free resolution of a finitely generated $A$-module $M$. We use the notation $\operatorname{reg}(M)$ when the ring $A$ is evident.
@@
[18, Lemma 3.1] Let $0\rightarrow M\rightarrow N\rightarrow P\rightarrow 0$ be a short exact sequence of finitely generated graded $A$-modules. Then, the following holds.
@@
The following are some properties of $d$-sequences and regular sequences that will be used in this article.
@@
    Let $\{f_{1},\ldots,f_{n}\}$ be a $d$-sequence in a ring $A$. Then, $(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n}^{s})=(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n})$ for all $s\geq 1$.
  2. (2)
    Let $\{f_{1},\ldots,f_{n}\}$ be a $d$-sequence in a ring $A$, and let $I=\langle f_{1},\ldots,f_{n}\rangle$. Then, for $s\geq 1$,
    ${(\langle f_{1},\ldots,f_{i-1}\rangle+I^{s}:f_{i})}=(\langle f_{1},\ldots,f_{i-1}\rangle:f_{i})+I^{s-1}$
    for all $1\leq i\leq n$ ([31, Observation 2.4]).
@@
The following are some results related to the Koszulness and Cohen-Macaulayness of the diagonals of bigraded algebras which will be used in the paper.
@@
    For $\Delta=(c,e)$ and a bigraded $K$-algebra $R$ with the $i^{th}$ Betti number corresponding to the shift $(a,b)$ being denoted by $\beta_{i,a,b}$, $R_{\Delta}$ is Koszul if $\max\{\dfrac{a}{c},\dfrac{b}{e}:\beta_{i,a,b}\neq 0\}\leq i+1$ for every $i$ ([10, Theorem 6.2]).
@@
Proposition 1.5.
[22, Proposition 4.10] Let $I$ be a homogeneous almost complete intersection in a polynomial ring over an infinite field, with the height of $I$ denoted by $n$. Then, $I$ is generated by a homogeneous $d$-sequence $\{f_{1},\ldots,f_{n+1}\}$ such that $\{f_{1},\ldots,f_{n}\}$ is a regular sequence.
@@ 2. Regularity of powers of Almost Complete Intersections
In this section, we extend their result to give the regularity of powers of almost complete intersections in the non-equigenerated case as well. Our result coincides with [31, Theorem 3.1] when the ideal is equigenerated. Note that an upper bound for the same is discussed in [21, Corollary 2.11].
@@
Consider the notation and assumptions as in Setup 1. Then, for $i=0,\ldots,n-1$,
@@
Since $J$ is generated by a regular sequence $\{f_{1},\ldots,f_{n-1}\}$ of degrees $d_{1},\ldots,d_{n-1}$, respectively, the Koszul complex corresponding to the sequence $\{f_{1},\ldots,f_{n-1}\}$ resolves the ideal $J$, and it has the following form:
$0\longrightarrow A(-(d_{1}+d_{2}+\ldots+d_{n-1}))\stackrel{{\scriptstyle\phi_{n-1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\phi_{k+1}}}{{\longrightarrow}}\bigoplus_{{1\leq i_{1}\leq i_{2}\leq\cdots\leq i_{k}\leq n-1}}A(-(d_{i_{1}}+d_{i_{2}}+\ldots+d_{i_{k}}))\stackrel{{\scriptstyle\phi_{k}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\bigoplus_{j=1}^{n-1}A(-d_{j})\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}A\longrightarrow 0,$ (1)
where $\phi_{i}^{\prime}s$ are the differentials in the Koszul complex.
[…]
$\operatorname{reg}\left(\dfrac{A}{J}\right)=\sum_{l=1}^{n-1}d_{l}-(n-1),$
which is greater than $\operatorname{reg}\left(\dfrac{A}{I}\right)$ by assumption. Hence,
@@
As a consequence of Note 1.3(1), we have $\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}\right)=\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}\rangle}\right)=\sum_{l=1}^{n-1}d_{l}-n+1-d_{n}.$ Then,
$\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}(-2d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{J}\right)+d_{n}$.
@@
Since $\{f_{1},\ldots,f_{n}\}$ forms a $d$-sequence with $\{f_{1},\ldots,f_{n-1}\}$ forming a regular sequence, from Note 1.3(2), we have $\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}):f_{i}}\right)(-d_{i})=\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i-1}\rangle:f_{i})+I}\right)(-d_{i}).$ From Note 1.3(3), this is equal to $\operatorname{reg}\left(\dfrac{A}{I}\right)+d_{i}<\sum_{l=1}^{n-1}d_{l}-(n-1)+d_{i}\leq\sum_{l=1}^{n}d_{l}-(n-1)\leq\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\right)+1$.
Hence, from Lemma 1.1(1),
@@
From Note 1.3(1), one obtains $\operatorname{reg}\left(\dfrac{A}{J:f_{n}^{s+1}}\right)(-sd_{n})=\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)+sd_{n}$. Then from Lemma 1.1(3),
@@
From Proposition 1.5, it is known that homogeneous almost complete intersection ideals in a polynomial ring over an infinite field have a generating set of $d$-sequence such that removing the last generator results in a regular sequence. Hence, as a consequence of Lemmas 2.1 and 2.2, one can determine the regularity of powers of almost complete intersections that satisfy certain conditions.
@@
Let $A$ be a standard graded polynomial ring over an infinite field $K$ and $I$ be a homogeneous almost complete intersection ideal of $A$ generated by a $d$-sequence $\{f_{1},\ldots,f_{n}\}$ in degrees $d_{1},\ldots,d_{n}$, where $\{f_{1},\ldots,f_{n-1}\}$ is a regular sequence and $J=\langle f_{1},\ldots,f_{n-1}\rangle$. Assume that $\operatorname{reg}(A/I)<\sum_{l=1}^{n-1}d_{l}-n+1$ and that $d_{n}=\max\{d_{i}\mid i=1,\ldots,n\}$ Then for all $s\geq 2$ and $i=0,\ldots,n-1$, the following is true,
@@
We prove the result by descending induction on $i$. Consider the base case of $i=n-1$ of the descending induction in Equation (4). Then clearly, $\langle f_{1},\ldots,f_{(n-1)+1}\rangle+I^{s}=I$. From Note 1.3(2) and Lemma 2.2,
$\operatorname{reg}\left(\dfrac{A}{(J+I^{s}):f_{n}}(-d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+I^{s-1}}\right)+d_{n}=\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+\langle f_{n}^{s-1}\rangle}\right)+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}\geq\sum_{l=1}^{n-1}d_{l}-(n-1)+1>\operatorname{reg}\left(\dfrac{A}{I}\right)+1$. Hence $\operatorname{reg}\left(\dfrac{A}{J+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}$ by Lemma 1.1(1).
Now assume $i\leq n-2$ and that the result holds for $i+1$. In the short exact sequence (4), by induction hypothesis on $i$,
$\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}.$
Since $\{f_{1},\ldots,f_{j}\}$ forms a $d$-sequence for $1\leq j\leq n$ and a regular sequence for $11\leq j\leq n-$ 1, from Note 1.3(2) and Note 1.3(3), $\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)=\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{s-1}}\right)$. This, by induction hypothesis on $s$, is equal to $\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}$.
Now, the final result is a consequence of the inequality,
$\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)+d_{i+1}\leq\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}+d_{n}<\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)+1$
and Lemma 1.1(1). ∎
@@
The above theorem fails if $d_{n}$ is not the maximal degree generator of $I$. For example consider $I=\langle f_{1},f_{2},f_{3},f_{4}\rangle=\langle x_{26}x_{45}-x_{25}x_{46}+x_{24}x_{56},x_{16}x_{45}-x_{15}x_{46}+x_{14}x_{56},x_{16}x_{25}x_{34}-x_{15}x_{26}x_{34}-x_{16}x_{24}x_{35}+x_{14}x_{26}x_{35}+x_{15}x_{24}x_{36}-x_{14}x_{25}x_{36},x_{36}x_{45}-x_{35}x_{46}+x_{34}x_{56}\rangle\subset K[x_{14},x_{15},x_{16},x_{24},x_{25},x_{26},x_{34},x_{35},x_{36},x_{45},x_{46},x_{56}]=B$.
Then $\{f_{1},f_{2},f_{3},f_{4}\}$ forms a $d$-sequence with $\{f_{1},f_{2},f_{3}\}$ being a regular sequence. But $\operatorname{reg}\left(\dfrac{B}{I^{3}}\right)=8$ which is not equal to $(2+2+2+3)-4+2=7$.
Note 2.5.
It is important to note that the formula for the regularity of powers in Theorem 2.3 holds true for a general setup of a homogeneous ideal $I$ in a standard graded $K$-algebra $A$, where the field is not necessarily infinite, provided that the following properties are satisfied:
@@
    $I$ has a generating set of $d$-sequence $\{f_{1},\ldots,f_{n}\}$ such that $\{f_{1},\ldots,f_{n-1}\}$ is a regular sequence.
@@
    $\operatorname{reg}(A/I)<\sum_{l=1}^{n-1}d_{l}-n+1$ and $d_{n}=\max\{d_{i}\mid i=1,\ldots,n\}$.
As a consequence of Theorem 2.3 and Note 2.5, the following can be observed regarding the regularity of powers of almost complete intersections of grade $3$. To understand the set of generators of these ideals, we look at the notion of Pfaffians.
Given a skew-symmetric matrix $X$, the Pfaffian of $X$ is defined as the square root of the determinant of $X$. For $l\in\mathbb{N}$, let $\operatorname{Pf}_{\bar{l}}(X)$ denote the Pfaffian of $X$ obtained by removing the $l^{th}$ row and the corresponding column of $X$, which can be seen as an element in $B=K[X]$. Let $X$ be skew-symmetric matrix of order $t$ with the following form:
$X=\begin{bmatrix}0&0&0&x_{14}&x_{15}&\cdots&x_{1t}\\ 0&0&0&x_{24}&x_{25}&\cdots&x_{2t}\\ 0&0&0&x_{34}&x_{35}&\cdots&x_{3t}\\ -x_{14}&-x_{24}&-x_{34}&0&x_{45}&\cdots&x_{4t}\\ -x_{15}&-x_{25}&-x_{35}&-x_{45}&0&\cdots&x_{5t}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ -x_{1t}&-x_{2t}&-x_{3t}&-x_{4t}&-x_{5t}&\cdots&0\\ \end{bmatrix}$ (5)
Then, from the results and discussions in [6] and [8], it can be concluded that an almost complete intersection ideal $I$ in $K[X]$ of grade $3$ and type at least $2$ arises as an ideal generated by Pfaffians of the following form,
  1. (1)
    If $t\geq 5$ is an odd integer, then
    $I=\langle\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X),\operatorname{Pf}_{\overline{123}}(X)\rangle.$ (6)
  2. (2)
    If $t\geq 6$ is an even integer, then
    $I=\langle\operatorname{Pf}(X),\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X)\rangle.$ (7)
Setup 2.
Let $B=K[X]$ denote a standard graded polynomial ring with $X$ being a skew-symmetric matrix of the form (5). For $t\in\mathbb{N}$ with $t\geq 5$, let $I$ be an almost complete intersection ideal of $B$ of grade $3$ and type $t-3$ (Equations (6) and (7)).
Remark 2.6.
Note that from [8, Theorems 1.2 and 1.3], one can obtain the minimal free resolutions of ideals of the form in Setup 2. As a consequence, the following properties can be observed.
  1. (1)
    $B/I$ is Cohen-Macaulay. Since $I$ is perfect in $B$, the observation follows from [13, Section 19].
  2. (2)
    For $t\geq 5$ and $I$ an ideal of type $t-3$, $\text{reg}(B/I)=t-4$.
  3. (3)
    For $t\geq 5$ and $I$ an ideal of type $t-3$, $(B/I)^{(c)}$ is Koszul for $c\geq\frac{t}{3}$.
@@
Let $X$ be a skew-symmetric matrix of order $t$ of the form given in Equation (5). Then,